OpenStudy (anonymous):

Fill in the missing exponents (5^4*5^2)/(5^3*5^7)^6=1/5 (?)

OpenStudy (anonymous):

(4 + 2) - (3 + 7) * (6) = -x, solve for x

OpenStudy (anonymous):

(5^4*5^2)/(5^3*5^7)^6 = 1/ (5^4)

OpenStudy (anonymous):

(5^4*5^2)/(5^3*5^7)^6 = 1/625 = 1/ (5^4) in the question : (5^4*5^2)/(5^3*5^7)^6 = 1/5 so, the missing exponent is 1/ (5^4)